• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name W. Deconinck

    Log entry time 19:08:37 on October25,2009

    Entry number 296661

    keyword=Reset the level meters, back to Compton studies

    I reset the level meters that were reporting bogus values in both the left and right arm. At this point we are getting 100% on Dipole Right (instead of 2% before a power cycle), so that should stabilize downwards. The flow meter A on Q1 R is still showing -16.0 l/m (I didn't touch it). MCC will keep an eye on our flows and report back if still problems in 30 mins. Until then back to beam permit and Compton studies.

    FIGURE 1